# The Distribution of Planet Radius in Kepler Multiplanet Systems Depends on Gap Complexity

### v1

@@ Abstract
The distribution of small planet radius ($<$4 R) is an indicator of the underlying processes governing planet formation and evolution. We investigate the correlation between the radius distribution of exoplanets in Kepler multiplanet systems and the system-level complexity in orbital period spacing. Utilizing a sample of 234 planetary systems with three or more candidate planets orbiting FGK main-sequence stars, we measure the gap complexity ($C$) to characterize the regularity of planetary spacing and compare it with other measures of period spacing and spacing uniformity. We find that systems with higher gap complexity exhibit a distinct radius distribution compared to systems with lower gap complexity. Specifically, we find that the radius valley, which separates super-Earths and sub-Neptunes, is more pronounced in systems with lower gap complexity ($C$$<$0.165). Planets in high complexity systems ($C$$>$0.35) exhibit a lower frequency of sub-Earths (2.5 times less) and sub-Neptunes (1.3 times less) and a higher frequency of super-Earths (1.4 times more) than planets in low complexity systems. This may suggest that planetary systems with more irregular spacings are more likely to undergo dynamic interactions that influence planet scattering, composition, and atmospheric retention. The gap complexity metric proves to be a valuable tool in linking the physical characteristics of planets to their orbital configurations.
@@ 1 Introduction
Around stars like the Sun on orbits shorter than than one year, the most common size of planet is between the radius of Earth and the radius of Uranus (1-4 R). Borucki et al. (2011) and Batalha et al. (2013) first identified this trend in planets observed by the Kepler space telescope. Occurrence rate studies (e.g. Howard et al. 2012; Petigura et al. 2013) which correct for the biases of observing transiting planets further confirm the overabundance of planets larger than the Earth but smaller than a gas-rich planet (5-10% gas by mass).
@@
From just observing planet radii, constraints can be placed on the formation and evolution of planets between 1-4 R. Small planets may either form from a bimodal amount of water (Luque & Pallé, 2022; Burn et al., 2024) or planets in the radius valley must undergo the loss of their volatiles (Owen & Wu, 2013). In the later case, planets above the valley are massive enough to hold onto light gasses, while planets in the valley lose gasses when internally heated (Ginzburg et al., 2018; Gupta & Schlichting, 2019, 2020) and when stripped by high energy photons (Owen & Wu, 2017; Van Eylen et al., 2018; Rogers et al., 2021).
@@
The “peas in a pod” observations constrain potential models for the formation of planetary systems. Plausible models to produce the intra-system uniformity in period spacing include the redistribution of planets via dynamical instability (Goldberg & Batygin, 2022; Lammers et al., 2023; Ghosh & Chatterjee, 2024), migration within the protoplanetary disk (Batygin et al., 2023; Goldberg & Batygin, 2023; Choksi & Chiang, 2023), and the minimization of energy within a low-mass system of planets (Adams et al., 2020).
@@
Can we link the radius distribution to properties of multis? Does the radius valley persist in the most regularly spaced multis verses multis with irregular spacings? In this work we will use gap complexity, detailed in Sect. 4, which only depends on planet period to summarize the spacing of planets in a system. If both the complexity and planet radius frequency change in tandem with one another we may be able to tie together formation models.
@@
Our paper is laid out as follows. In the next section, Sec. 2, we describe and review the literature on gap complexity. In Sect. 3, we define our sample of planet candidates in multis around FGK main-sequence stars observed by Kepler. We find the gap complexity of these systems and report any correlations with underlying variables in Sec. 4. In Sec. 5, we show how the planetary radius distribution changes in systems of increasing gap complexity and changes when splitting systems into low and high gap complexity (Sec. 5.1). We discuss changes in the radius distribution in Sec. 6 and conclude in Sec. 7.
2 Gap Complexity: A Primer
@@
How far planets are spaced from one another in a planetary system is a strong control on the stability time of the planets’ orbits and the timescale of their dynamical evolution (Fang & Margot, 2013; Chambers et al., 1996; Obertas et al., 2017; Rice et al., 2018). However, for dynamical considerations spacing is generally measured in multiples of mutual Hill Radii which depends on planet mass. In this work with Kepler planets, we have well determined periods but limited mass measurements. Additionally, while many Kepler systems are near commensurability with mean motion resonance, the majority are not in resonance (Fabrycky et al., 2014; Steffen & Hwang, 2015), so we do not measure planet spacing based upon distance to a mean motion resonance. Thus, we adopt the gap complexity as a measure to describe the architecture of multis.
@@
[…]
3 System Selection
The recent catalog of planet candidates observed by the Kepler space telescope, Lissauer et al. (2023), found 253 planetary systems of three or more candidate planets. One focus of this catalogue which contains 4376 candidates and 709 multis is providing accurate planet parameters in multiplanet systems. The radius distribution is likely different around low-mass stars (Morton & Swift, 2014; Van Eylen et al., 2021; Ho et al., 2024) and the catalogue has few evolved stars. Thus, we limit systems to FGK main-sequence stars between 4000 and 7000 K which removes 12 systems and with surface gravity over 104 which removes an additional 3 systems. These cuts are similar to those in Hsu et al. (2018).
With these cuts there are 828 candidate planets in 238 planetary systems which have three or more planets. The periods and radii of the planets are shown in Fig. 1. The maximum error in the planet period is 0.04% with a median error below 10-3%.
In this work we aim to study the gap complexity in the inner planets of a system. Thus, we limit the period to 1000 days which removes three of the longest period planets. Other works make more stringent cuts. He & Weiss (2023) limits systems to one year periods. This limit would remove 15 planets and would result in three systems no longer having three planets. The completeness of the Kepler survey begins to drop at periods of 100 days. A 100 day cut removes 12% of planets and results in 16% of systems (38) no longer having 3 planets.
In order to use the most complete systems possible we only cut periods at 1000 days. In Sect. 4, we show the gap complexity with the three period cuts discussed above. For the radius distribution reported later in this work, we only show results for the 1000 day period cut. However, we do find the results hold for a 100 day cut with only slightly less significance primarily from the number of systems decreasing by 16%.
4 Gap Complexity of Kepler Multis
@@
In Fig. 2, we show an empirical cumulative distribution function of gap complexity for our systems. We show three lines with the three maximum period cut-offs. The most stringent cut is at most 7% different then our relaxed 1000 day cut. The planets near 1000 days move a few low complexity systems to the highest complexities visible as a slight uptick in systems above $C$=0.9. The distribution is in agreement with the DR25 catalogue of Kepler candidate planets (also analyzed in He & Weiss (2023)) differing by at most 2.5% between $C$=0.5 and 0.75.
Figure 2: The gap complexity for Kepler systems of three or more planets with a various outer planet period cut-off. We use the maximum 1000 day cutoff in this work. Over 50% of systems have a gap complexity under the 0.126 of the inner Solar System.
@@
A majority of Kepler multis have gap complexities under that of the inner Solar System (0.126). Half of systems are below $C$=0.1057 and 13% of systems are below $C$=0.01. While the 10% of systems with highest $C$ spread between $C$=0.66 and the maximum of 0.986 in the Kepler-123 3-planet system.
A majority of our 238 systems are 3-planet systems making up 68.5% of the systems. 20.2% are in 4-planet systems. 11.3% are in higher multiplicity systems (21 in 5-planet, 4 in 6-planet, 1 in 7-planet, and 1 in 8-planet systems).
4.1 Correlations with Other Measures of Period
@@
We find that stellar parameters have no correlation with gap complexity. The sample Pearson correlation coefficients, $r$, between $C$ and stellar temperature, mass, radius, log(g), and metallicity are all below 0.1 with slopes consistent with zero. Additionally, the average radius uncertainty in a system and the average log-flux in a system has no correlation with gap complexity.
@@
However, the gap complexity is sensitive to other period-related measures. While not correlated with inner period, the outer period has a correlation coefficient of 0.37 with $C$. From a slope, $b$, of a linear regression ($y=a+bx$), the outer planet period around a $C$=1.0 system is expected to be 155 days longer than a $C$=0 system. This correlation still exists if the maximum period allowed for a system is set to 100 days (the slope of outer period with $C$ is then 20 days).
@@
Correlations also exist between gap complexity and the range of periods in log-space ($b=0.72$ days, $r=0.47$) and the normalized standard deviation of period ($\sigma_{P}/\mu_{P}$, $b=0.53$, $r=0.49$). The strongest correlation, $r=0.64$, is shown in Fig. 3 top between gap complexity and the average $\Delta P/P_{i}$ in a system. Where the average $\Delta P/P_{i}$ is equal to
@@
Since the ratio of the outer planet to inner planet period in each nearest pair is equal to $1+\Delta P/P_{i}$, the figure shows a majority of systems with C$<$0.1 have average period ratios less than 3:1. While systems at large gap complexities may have low period ratios, they have a wider spread in period ratios with some over 10:1. These correlations remain significant if we limit periods to 100 days with slightly lower slopes and correlation coefficients.
Figure 3: Top, the average $\Delta P/P_{i}$ in the systems against their gap complexities. Low-$C$ systems have planets close together in period while the dispersion in period spacing is much larger in high-$C$ systems. Bottom, the Gini index of planet period ratio verses gap complexity for our systems. These two measures of period spacing uniformity are highly correlated. A linear regression fit is shown with the slope and correlation coefficient ($r$) reported.
We find that systems with larger gap complexities tend to be more widely spread out in planet periods. This has not been considered in previous works and should be considered as an underlying variable in the subsequent sections.
Additionally, other works use different measures to describe how planets are spaced in a system. Following upon the work of Goyal & Wang (2022) and Goyal et al. (2023), Goyal & Wang (2024) describes the spacing uniformity in systems with the adjusted Gini index (Deltas, 2003; Gini, 1912) of a system’s period ratios ($\mathcal{G}_{P_{ratio}}$). The Gini index is an economic measure of inequality and is used here to describe the uniformity of period ratios in a system. In Fig. 3 bottom, we show the correlation between gap complexity and the Gini index of period ratios. The Gini index is also normalized to be between zero and one, but the maximum in our systems is 0.88. While the correlation is not 1:1 the two parameters to measure the uniformity of spacing in systems are highly correlated.
@@ 5 Changes in the Radius Distribution with Gap Complexity
We calculate gap complexity using all the planets in our system sample, but now we limit our radius analysis to planets under 6 R and with radius uncertainties below 50%. This limits the analysis to 783 planets in 234 systems. Four systems go from having three planets to no planets which pass this criteria. Fourteen systems have less than three planets, 151 systems have three planets, and 69 systems have more than three planets. The gap complexity of the four systems with no planets which pass our criteria are $C$=0.08, 0.13, 0.26, 0.34.
@@
In the subsequent section, we will split our sample into low- and high-$C$ to examine the radius distribution in detail. As a first test, we want to find the significance of the difference in radius frequency at higher/lower gap complexities than the median. To do so, we design the following test of the running average of $C$ with planet radius. In Fig. 5, we show the running median $C$ in 80 planets ordered by their observed median radius. We then create test samples where we shuffle the gap complexities and randomly assign them to each system. The median $C$ across all radii for 500 of these test samples where radius and gap complexity is not correlated is 0.22$\pm$0.05.
@@ 5.1 Radii in Systems of low- and high-C
In this section, we aim to examine the radius distribution in systems of low- and high-$C$ in more detail. First, we determine at what point to divide the systems. We perform a two-sample Kolmogorov–Smirnov (K-S) test between samples of planet radius divided into low-/high-$C$ with various maximum/minimum values of $C$ determining the samples. A two-sample K-S test gives the probability that both samples were drawn from the same underlying population without assumption of that underlying distribution. We emphasize that the K-S test is sensitive to sample size and cannot be used to tell how different or in what ways two distributions are different. For example, a p-value less than 0.05 can be achieved by drawing 30 values from $\mathcal{N}(0,1)$ and from $\mathcal{N}(1,1)$ or similarly by drawing 3000 values from $\mathcal{N}(0,1)$ and from $\mathcal{N}(0,0.9)$.
@@
Figure 6: We perform K-S tests between the radius of planets separated by the system’s gap complexities ($C$). One radius sample has a maximum $C$ given by the x-axis and the other has a minimum set by the y-axis. The significance when using all planets from 0-6 R is shown with purple x’s. While the significance for planets between 1-3 R is shown with red o’s. Size and darkness of the marker represents the significance value without considering radius uncertainties.
In Fig. 6, we show the p-value of the K-S test with different divisions. When using all planet radii, the most significance occurs when the sample is divided into systems with $C$$<$0.105 and systems of $C$$>$0.11. We see the p-value drops quickly when the maximum $C$=0.105-0.11 while the p-value is less dependent on the minimum $C$.
However, our primarily interest is in in the region between one and three Earth-radii which features the radius valley. Thus, we also perform K-S tests between two samples of planets in this region in systems divided by gap complexities. The significance comes at a higher maximum gap complexity of $C$$<$0.165 and a much higher minimum gap complexity of $C$$>$0.35. We see the significance fall off in all directions from this point relatively equally. Above a $C$$>$0.3 there is significance across all cuts of low gap complexity.
@@
In Fig. 7 left, we show the radius distribution for systems below and above the determined gap complexities. There are 118 systems with $C$$<$0.105 and 117 systems with $C$$>$0.11 and 3 systems between. We see when considering all planets the abundance of sub-Earths at low-$C$ dominate the significance. For each of the low- and high-$C$ systems, we show 100 distributions where each planet’s radius is randomly chosen from a Gaussian with observed median value and a standard deviation equal to the average of the plus/minus observed uncertainties. We calculate the p-value of a 2-sample K-S test between each sampled radius distribution for the 100 distributions. The median p-value of a 2-sample K-S test between these two distributions is 0.0044.
@@
For the region of 1-3 R, we also perform a test of modality on the low- and high-$C$ planets. Using a Hartigan dip-test of unimodality at a p-value of 0.01 and 1000 trials we find the low-$C$ planets to have two peaks while the high-$C$ only has one peak. The test find the two peaks to be from 1.08 to 1.9 R referred to as super-Earths and from 1.90 to 2.40 R referred to as sub-Neptunes in this work. While for the high-gap complexity planets, the test returns one broad peak from 1.01 to 2.98 R.

### v2

@@ Abstract
The distribution of small planet radius ($<$4 R) is an indicator of the underlying processes governing planet formation and evolution. We investigate the correlation between the radius distribution of exoplanets in Kepler multiplanet systems and the system-level complexity in orbital period spacing. Utilizing a sample of 234 planetary systems with three or more candidate planets orbiting FGK main-sequence stars, we measure the gap complexity ($C$) to characterize the regularity of planetary spacing and compare it with other measures of period spacing and spacing uniformity. We find that systems with higher gap complexity exhibit a distinct radius distribution compared to systems with lower gap complexity. Specifically, we find that the radius valley, which separates super-Earths and sub-Neptunes, is more pronounced in systems with lower gap complexity ($C$$<$0.165). Planets in high complexity systems ($C$$>$0.35) exhibit a lower frequency of sub-Earths (2.5 times less) and sub-Neptunes (1.3 times less) and a higher frequency of super-Earths (1.4 times more) than planets in low complexity systems. This may suggest that planetary systems with more irregular spacings are more likely to undergo dynamic interactions that influence planet scattering, composition, and atmospheric retention. The gap complexity metric proves to be a valuable tool in linking the orbital configurations of planets to their physical characteristics.
@@ 1 Introduction
Around stars like the Sun on orbits shorter than than one year, the most common size of planet is between the radius of Earth and the radius of Uranus (1-4 R). Borucki et al. (2011) and Batalha et al. (2013) first identified this trend in planets observed by the Kepler space telescope. Occurrence rate studies (e.g. Howard et al. 2012; Petigura et al. 2013; Burke et al. 2015) which correct for the biases of observing transiting planets further confirm the overabundance of planets larger than the Earth but smaller than a gas-rich planet (5-10% gas by mass).
@@
From just observing planet radii, constraints can be placed on the formation and evolution of planets between 1-4 R. Small planets may either form from a bimodal amount of water (Luque & Pallé, 2022; Burn et al., 2024) or planets in the radius valley must undergo the loss of their volatiles (Owen & Wu, 2013; Chen & Rogers, 2016). In the later case, planets above the valley are massive enough to hold onto light gasses, while planets in the valley lose gasses when stripped by high energy photons (Owen & Wu, 2017; Van Eylen et al., 2018; Rogers et al., 2021) and when internally heated by both core formation (Ginzburg et al., 2018; Gupta & Schlichting, 2019, 2020) and giant impacts (Biersteker & Schlichting, 2019).
@@
The “peas in a pod” observations constrain potential models for the formation of planetary systems. Plausible models to produce the intra-system uniformity in period spacing include the redistribution of planets via dynamical instability (Goldberg & Batygin, 2022; Lammers et al., 2023; Ghosh & Chatterjee, 2024), migration within the protoplanetary disk (Batygin et al., 2023; Goldberg & Batygin, 2023; Choksi & Chiang, 2023; Shariat et al., 2024), and the minimization of energy within a low-mass system of planets (Adams et al., 2020).
@@
The mechanism that lead to the regular spacing of planets in multis may also affect the composition of the planets formed. In this work, we examine the frequency of planets near the radius valley in multis with uniform and irregular spacings. To summarize the spacing of planets in multis, we calculate a system’s gap complexity which only depends on planet period. If both the complexity and planet radius frequency change in tandem with one another we may be able to tie together formation models.
[…]
@@
Our paper is laid out as follows. In Sect. 2, we define our sample of planet candidates in multis around FGK main-sequence stars observed by Kepler. We find the gap complexity of these systems and report any correlations with underlying variables in Sec. 3. In Sec. 4, we show how the planetary radius distribution changes in systems of increasing gap complexity and changes when splitting systems into low and high gap complexity (Sec. 4.1). We discuss changes in the radius distribution in Sec. 5 and conclude in Sec. 6.
2 System Selection
The recent catalog of planet candidates observed by the Kepler space telescope, Lissauer et al. (2023), found 253 planetary systems of three or more candidate planets. One focus of this catalogue which contains 4376 candidates and 709 multis is providing accurate planet parameters in multis. The radius distribution is likely different around low-mass stars (Morton & Swift, 2014; Van Eylen et al., 2021; Ho et al., 2024) and the catalogue has few evolved stars. Thus, we limit systems to FGK main-sequence stars between 4000 and 7000 K which removes 12 systems and with surface gravity over 104 which removes an additional three systems. These cuts are similar to those in Hsu et al. (2018).
In this work we aim to study the gap complexity in the inner planets of a system. Thus, we remove planets with orbital periods over 1000 days. This removes three of the longest period planets while keeping the systems as complete as possible for the calculation of gap complexity.
Our sample of multis includes 825 candidate planets in 238 planetary systems which have three or more planets. The periods and radii of the planets are shown in Fig. 1. The maximum error in the planet period is 0.04% with a median error below 0.001%. A majority of our 238 systems are 3-planet systems making up 68.5% of the systems. 20.2% are in 4-planet systems. 11.3% are in higher multiplicity systems (21 in 5-planet, 4 in 6-planet, 1 in 7-planet, and 1 in 8-planet systems).
Other works have more stringent limits on the maximum orbital period. He & Weiss (2023) limits systems to one year periods. This limit would remove 15 planets and would result in three systems no longer having three planets. The completeness of the Kepler survey begins to drop at periods of 100 days. A 100 day cut removes 12% of planets and results in 16% of systems (38) no longer having three planets compared to our 1000 day cut. While not presented in this work, our findings on the changes in the radius distribution with gap complexity (Sec. 4) hold for a 100 day period cutoff with only slightly less significance primarily from the number of systems decreasing by 16%.
3 Gap Complexity of Kepler Multis
@@
The stability time of a planetary system strongly depends on how far planets are spaced from one another within the system (Fang & Margot, 2013; Chambers et al., 1996; Obertas et al., 2017; Rice et al., 2018). However, for dynamical considerations spacing is generally measured in multiples of mutual Hill Radii which depends on planet mass. In this work with Kepler planets, we have well determined periods but limited mass measurements. Additionally, while many Kepler systems are near commensurability with mean motion resonance, the majority are not in resonance (Fabrycky et al., 2014; Steffen & Hwang, 2015), so we do not measure planet spacing based upon distance to a mean motion resonance. Thus, we adopt the gap complexity as a measure to describe the architecture of multis.
@@
In Fig. 2, we plot the orbital architectures of 20% of the systems in our sample ordered by their gap complexity. The systems are chosen so that the proportions of systems below/above a given $C$ is representative of the entire sample. As $C$ increases the spacing between planets in log-period space becomes visually more diverse.
@@
Additionally in Fig. 3, we show an empirical cumulative distribution function of gap complexity for our systems. The distribution is in agreement with the DR25 catalogue of Kepler candidate planets (NASA Exoplanet Archive 2024, also analyzed in He & Weiss (2023)). With our relaxed orbital period cutoff of 1000 days a few systems with low complexity in shorter period planets move to the highest complexities when the exterior planet is included. This is visible as a slight uptick in systems with $C$$>$0.9.
@@
A majority of Kepler multis have gap complexities under that of the inner Solar System (0.126). Half of systems are below $C$=0.1057 and 13% of systems are below $C$=0.01. While the 10% of systems with highest $C$ spread between $C$=0.66 and the maximum of 0.986 in the Kepler-123 (KOI-238) 3-planet system.
Figure 2: Architectures of planetary systems where each horizontal line is a system with the star’s KOI number and the system’s $C$ noted. Every fifth planetary system when the 238 systems are ordered by $C$ (48 systems) are shown. Points are placed for each planet at its orbital period with a size proportional to the planet radius. Planets are colored by size category. Unfilled planets are over 6 R or have radius uncertainties over 50%. Unfilled planets are used to calculate the gap complexity, but the radius is not used in later analyses. The inner Solar System is also shown.
3.1 Correlations with Other System Properties
@@
We find that the gap complexity of a system does not correlate with the host star’s properties. The sample Pearson correlation coefficients, $r$, between $C$ and stellar temperature, mass, radius, log(g), and metallicity are all below 0.1 with slopes consistent with zero. The average radius uncertainty and the average log-flux of planets in a system has no correlation with gap complexity. Additionally, we find sensitivity-related parameters such as the Multiple Event Statistic (MES) have no correlation with gap complexity both when all values for the planets are used and when values are averaged in a system. We also find the distributions of MES in low and high complexity systems are consistent with each other.
@@
However, the gap complexity is sensitive to other period-related measures. While not correlated with inner period, the outer period has a correlation coefficient of 0.37 with $C$. From a slope, $b$, of a linear regression ($y=a+bx$), the outer planet period around a $C$=1.0 system is expected to be 155 days longer than a $C$=0 system.
@@
Correlations also exist between gap complexity and the range of periods in log-space ($b=0.72$ days, $r=0.47$) and the normalized standard deviation of period ($\sigma_{P}/\mu_{P}$, $b=0.53$, $r=0.49$). The strongest correlation, $r=0.64$, is shown in Fig. 3 left axis between gap complexity and the average $\Delta P/P_{i}$ in a system. Where the average $\Delta P/P_{i}$ is equal to
@@
Since the ratio of the outer planet to inner planet period in each nearest pair is equal to $1+\Delta P/P_{i}$, the figure shows a majority of systems with C$<$0.1 have average period ratios less than 3:1. While systems at large gap complexities may have low period ratios, they have a wider spread in period ratios with some over 10:1.
Figure 3: Black, empirical cumulative distribution function of gap complexity for our 238 systems. Purple circles, the average $\Delta P/P_{i}$ in the systems (left axis) against their gap complexities. Low-$C$ systems have planets close together in period while the dispersion in period spacing is much larger in high-$C$ systems. Red diamonds, the Gini index of planet period ratio (right axis) verses gap complexity for our systems. These two measures of period spacing uniformity are highly correlated. For both the purple and red points a linear regression fit is shown with the slope and correlation coefficient ($r$) reported.
We find that systems with larger gap complexities tend to be more widely spread out in planet periods. This has not been considered in previous works and for our purposes is considered as an underlying variable in the subsequent sections.
Additionally, other works use different measures to describe how planets are spaced in a system. Following upon the work of Goyal & Wang (2022) and Goyal et al. (2023), Goyal & Wang (2024) describes the spacing uniformity in systems with the adjusted Gini index (Deltas, 2003; Gini, 1912) of a system’s period ratios ($\mathcal{G}_{P_{ratio}}$). The Gini index is an economic measure of inequality and is used here to describe the uniformity of period ratios in a system. In Fig. 3 right axis, we show the correlation between gap complexity and the Gini index of period ratios. The Gini index is also normalized to be between zero and one, but the maximum in our systems is 0.88. While the correlation is not 1:1 the two parameters to measure the uniformity of spacing in systems are highly correlated.
@@ 4 Changes in the Radius Distribution with Gap Complexity
We calculate gap complexity using all the planets in our system sample with orbital periods less than 1000 days, but now we limit our radius analysis to planets under 6 R and with radius uncertainties below 50%. This limits the analysis to 783 planets in 234 systems. Four systems go from having three planets to no planets which pass this criteria. Fourteen systems have less than three planets, 151 systems have three planets, and 69 systems have more than three planets. The gap complexity of the four systems with no planets which pass our criteria are $C$=0.08, 0.13, 0.26, 0.34.
@@
As a first test of the significance of these features, we test if a given radius prefers systems of higher or lower $C$ than the median $C$. In Fig. 5 the thick red line shows the running median $C$ of 80 planets when the planets are ordered by their observed median radius. We then create test samples where we randomly assign the gap complexities in our sample to each system. Each thin blue line in Fig. 5, is the running median $C$ in a test sample where the gap complexity has been randomly assigned. The median $C$ across all radii for 500 of these test samples is 0.22$\pm$0.05.
@@ 4.1 Radii in Systems of low- and high-C
In this section, we aim to examine the radius distribution in systems split into samples of low- and high-$C$. First, we determine at what gap complexity to divide the systems. We perform a two-sample Kolmogorov–Smirnov (K-S) test between samples of planet radius divided into low-/high-$C$ with various maximum/minimum values of $C$ determining the samples. A two-sample K-S test gives the probability that both samples were drawn from the same underlying population without assumption of that underlying distribution. We emphasize that the K-S test is sensitive to sample size and cannot be used to tell how different or in what ways two distributions are different. For example, a p-value less than 0.05 can be achieved by drawing 30 values from $\mathcal{N}(0,1)$ and from $\mathcal{N}(1,1)$ or similarly by drawing 3000 values from $\mathcal{N}(0,1)$ and from $\mathcal{N}(0,0.9)$.
@@
Figure 6: We perform K-S tests between the radius of planets separated by the system’s gap complexities ($C$). One radius sample has a maximum $C$ given by the x-axis and the other has a minimum set by the y-axis. The significance when using all planets from 0-6 R is shown with purple x’s. While the significance for planets between 1-3 R is shown with red circles. Size and darkness of the marker represents the significance value without considering radius uncertainties.
In Fig. 6, we show the p-value of the K-S test with different divisions. When using all planet radii ($<$6 R), the most significance occurs when the sample is divided into systems with $C$$<$0.105 and systems of $C$$>$0.11. However, our primarily interest is in in the region between one and three Earth-radii which features the radius valley. Thus, we also perform K-S tests only considering planets in this region. The significance comes when low-$C$ is defined as $C$$<$0.165 and high-$C$ is defined as $C$$>$0.35.
Moving forward we will divide out samples into low- and high-$C$ based upon the minimum p-value of the K-S tests since no division exists a priori. However, Fig. 6 shows these points are not exceptional with significance varying slowly and smoothly in most directions from the chosen division. When only considering the planets with 1$<$R$<$3 R, there is significance across all cuts of low-$C$ for any high-$C$ cut above 0.3.
@@
In Fig. 7, we show the radius distribution for systems below and above the determined gap complexities. For each of the low- and high-$C$ systems, we show 100 distributions where each planet’s radius is randomly chosen from a Gaussian with observed median value and a standard deviation equal to the average of the plus/minus observed uncertainties. We calculate the p-value of a 2-sample K-S test between each sampled radius distribution for the 100 distributions.
When considering all planets, there are 118 systems with $C$$<$0.105 and 117 systems with $C$$>$0.11 and 3 systems between. The median p-value of a 2-sample K-S test between these two distributions is 0.0044. We see when considering all planets, Fig. 7 left, the abundance of sub-Earths at low-$C$ dominate the significance.
@@
For the region of 1-3 R, we also perform a test of modality on the low- and high-$C$ planets. Using a Hartigan dip-test of unimodality at a p-value of 0.01 and 1000 trials we find the low-$C$ planets to have two peaks while the high-$C$ only has one peak. The test find the two peaks to be from 1.08 to 1.9 R and from 1.90 to 2.40 R. While for the high-gap complexity planets, the test returns one broad peak across the analysis range from 1.01 to 2.98 R.